Standard USB phone charges are 5 V. There are two common types of charger, 1 A and 2.1 A. A 1 A charger is capabable of delivering up to 1 A of current. A 2.1 A charged can deliver more. 1. What is the maximum power each charger can output? 2. If a phone is connected to a standard USB charger and draws 2.1 A of current, what is the phone’s resistance? 3. Some phone manufacturers make USB chargers that can produce more power than a standard USB charger. For example, Motorola makes a “TurboPower” charger that can produce 15W of power. It does this by increasing the voltage. For example, the TurboPower Charger can operate at 10 V. What is the maximum amount of current that can be drawn from a TurboPower charger operating at 10 V?
Answer #1

a)

Maximum Power each Charger can output is

P1A = VI =5*1=5 W

P2.1A=5*2.1 =10.5 W

b)

Phone's Resistance is

R=V/I =5/2.1 =2.381 ohms

c)

Current drawn by turbocharger is

I=P/V =15/10 = 1.5 A